OpenStudy (anonymous_user):

What is the correct solution to the equation? 4(x – 3) = 16 A. 1 B. 7 C. 15 D. 24 I think that it is B. 7 Am I right?

OpenStudy (anonymous):

B

OpenStudy (anonymous_user):

thanks!

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

dang lol i was about to say that but you answered faster

OpenStudy (anonymous):

4(x-3)=16 4x-12=16 4x=28 4x/4=28/4 x=7